Amanda is six times as old as Michael. In 16 years, Amanda will be only twice as old as Michael. What are their ages now?
I'm not sure where to even start with this problem. Please help. Thanks

Answer by stanbon(75887)

Amanda is six times as old as Michael. In 16 years, Amanda will be only twice as old as Michael. What are their ages now?
-----
Equations:
a = 6m
a+16 = 2(m+16)
-----
Substitute for "a" and solve for "m"::
6m + 16 = 2m + 32
----
4m = 16
m = 4 yrs (Michael's age now)
------
Solve for "a":
a = 6m = 6*4 = 24 yrs (Amanda's age now)
------------
Cheers,
Stan H.
--------------